Shift summary night 02 02 2000


Subject: Shift summary night 02 02 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:50:34 CST


Shift Summary for night shift, 02 02 2000
(the date is interesting, and no one predicted the end of
civilization for these number combinations)

On Shift: Nick and Bruce

Hours of beam delivered at all intensities: 7.5
Hours of beam delivered at nominal intensity: 5.5
Hours of KAMI data: 5.5

Accounting of lost time: targeting angle changed to 15 mrad

===============================================================
Runs ended this shift: 15741-15746

===============================================================
Tapes brought to FCC this shift: nonr

Accesses during shift: Bruce and Wah go on access so Bruce can
                       look around.
===============================================================
Description of Shift: Beam scans, then low intensity runs with
different Be thicknesses to get n/K0 ratio.

run spill time comment
----------------------------------------------------
             1530 hr Nick takes over shift, Wah and Doug call
                       appologizing that Bruce will be late.
             1615 hr Rick F from MCR calls they are ready to
                       deliver beam, we requested muons and
                       paged Doug J.
15741 1625 hr Muon run started with 15 mrad targeting
                       angle. As soon as accelertor gets to
                       high enought intensity Rick F. with call
                       then we will go to hadron mode and he will
                       do an hour or so of targeting angle studies.
             1626 hr Bruce arrives.
         9 1635 hr Beam goes away.
        14 Beam is back.
        17 Beam goes away, and varying intensity.
        21 1650 hr Trigger of muon run stopped for access.
        35 1705 hr Access ended, triggers started, back to
                       muons, MCR still tuning beamline, now
                       downstream of switchyard.
        86 1815 hr Muon run ends, openning beam stop going
                       to KAMI hadron.
             1820 hr Initlizing Kami hadron.
15742 1 1825 Run starts, now we start hardon beam scan.
             1845 hr Finished first beam scans, now doing
                       off of center.
             2000 hr !
             2002 hr !!
             2010 hr Finished three scans.
15742 84 2012 hr Run stopped, just a junk run with beam scans.
15743 1 2016 hr Run started, 0" Be, 5e11 intensity.
             2020 hr 2 and 0 in date and computers did not killed us
                       yet!
             2022 hr dito (humm)
15744 1 2035 hr Run started, 10" Be, 5e11 intensity.
15745 1 2115 hr Run started, 20" Be, 5e11 intensity.
15746 1 2202 hr Run started, 38" Be, 5e11 intensity.
             2222 hr Look at all those 2 in the date and time!
                       (now it is over no more hours with just 2 and 0
                       left tonight), this is a boring shift no
                       kfrends 2 do.
             2330 hr Bruce leaves, quite control room.
             2336 hr Run stopped.
                       Rick changes threshold [from 50 to 30 mV] on ucal
                       analog sum threshold for new run to investigate
                       strange peak at 38 " Be.
15747 1 2342 hr Run started.
          4 Strange peak is larger, conclusion by Rick is it
                       is not stright threw muons.
          9 Run stopped.
             2354 hr Called MCR, asked for a special request, Max
                       said "he would put on his DJ hat and spin any
                       special request we want" (this run will end soon).
                       We asked for magnets to be set to muon mode but
                       leave the beam stop open. Rick K put the threshold
                       on ucal back to 50 mV.
15748 1 2356 hr Run starts.
             2359 hr Max calls back and says that he does not like
                       the special request, concerned that it is not
                       written down.
             2400 hr Shift ends with Rick K on the phone with Max
                       arguing it out based on rates.

=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=-=
Dr. Nickolas Solomey, Research Scientist, Enrico Fermi Institute
The University of Chicago, 5640 S. Ellis Ave, Chicago, Ill. 60637 USA
Phone: (773) 702-3278, fax: x-6317, http://hep.uchicago.edu/~solomey



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:50:39 CST